If cos x = -1/5, find the value of cos(x+pi)
Answer by lwsshak3(11628)

If cos x = -1/5, find the value of cos(x+pi)
***
sinx=√(1-cos^2x)=√(1-1/25)=√(24/25)=√24/5
Identity: cos(a+b)=cosa*cosb-sina*sinb
..
cos(x+π)=cosx*cosπ-sinx*sinπ=(-1/5*-1)-(√24/5*0)=1/5